# Anisotropy of Nanohertz Gravitational Wave Background and Individual Sources from Supermassive Binary Black Holes Based on Cosmological Simulation

### v2

@@ 3 Results
With methods described in Section 2, we produced $N_{\rm total}\approx 3.027\times 10^{7}$ light cone SMBBH GW sources distributed in the observable universe. Each binary system has a distinct location, redshift, binary black hole masses, and gravitational wave frequency. We superimposed the gravitational wave signals from these binary black hole systems to obtain the GWB signal and studied its properties. In this section, we will present the basic characteristics of the SMBBH mock data we have obtained, as well as our main results for the GWB signal.
3.1 Statistical Properties of Light Cone SMBBH Mock Data
We present the basic statistical properties of our light cone SMBBH mock data in this subsection. The total mass $M$ (left panel) and mass ratio $q$ (right panel) distributions of the SMBBHs are demonstrated in Figure 3. The total masses of the binary systems are basically between $10^{6}M_{\odot}$ and $10^{9.5}M_{\odot}$ and peak around $5\times 10^{7}M_{\odot}$. On the other hand, the mass ratio distribution of the SMBBHs generally exhibits a monotonically decreasing trend with increasing mass ratio $q$, the vast majority of the binary black hole systems have mass ratios below 0.4, and even close to zero. This indicates that in our data, the majority of the binary black hole systems exhibit a significant difference in mass between the two black holes. The redshift distribution of the SMBBHs is shown in Figure 4. This distribution peaks around redshift $z\sim 1.1$, and the maximum redshift can reach $z=6.5$.
The frequency distribution of the GW signals from our SMBBH systems, including rest frame frequency $f_{\rm r}$ and observed frequency $f_{\rm obs}$, are shown in Figure 5. They basically follow the power law distribution as expected from Eq. (4). Since $f_{\rm obs}$ is the redshifted $f_{\rm r}$, so we usually have $f_{\rm obs}<f_{\rm r}$, we can see from the figure that the distribution range of $f_{\rm obs}$ is correspondingly lower.
Figure 3: The total mass $M$ (left) and mass ratio $q$ (right) distribution of the light cone SMBBHs.
Figure 4: The redshift distribution of the light cone SMBBHs. The distribution peak at around $z\sim 1.1$, while the maximum redshift is 6.2.
Figure 5: The frequency distribution of the GW sources, red: rest frame frequency $f_{\rm r}$; green: observed frequency $f_{\rm obs}$. They basically follow power law behavior as expected. The distribution range of $f_{\rm obs}<f_{\rm r}$ is relatively lower due to redshift.
@@ 3.2 Isotropic Part of the Simulated GWB Signal
Unlike a series of earlier studies which relied on the estimation of SMBBH merger rate to deduce the characteristic amplitude of the GWB signal, our methodology incorporates a vast amount of individual sources scattered across the observable universe to construct GWB. From the green curve in Figure 5 we can see that, The observable frequency range of our GWB signal is roughly between $10^{-10}$z to $10^{-6}$z. Considering $10^{-10}$z is already below the lower frequency limit of the PTA band. We will primarily analyze the behavior of the resulted GWB signal between $10^{-9}$z to $10^{-6}$z.
@@
To calculate the resulting monopole part of the GWB signal and its characteristic strain spectrum, we divide $\ln f$ into $N_{\rm bin}=10$ bins in the range $10^{-9}$z to $10^{-6}$z, and select those SMBBHs whose observed frequencies are within the specified range in each bin. The resulting GW characteristic strain near a certain frequency can be calculated as a summation of strains from selected individual GW sources in the respective bin divided by the bin width, which can be expressed as:
$h_{\rm c}(f_{j})=\sqrt{\sum_{\ln f_{j}-\frac{\Delta\ln f}{2}<\ln f_{i}\leq\ln f_{j}+\frac{\Delta\ln f}{2}}\frac{h_{i}^{2}}{\Delta\ln f}},$
@@
where $\Delta\ln f=\frac{f_{\max}-f_{\min}}{N_{\rm bin}}$, and $h_{i}$ can be computed according to Equation (1). We present our results regarding the characteristic strain spectrum of GWB in Figure 6. In this figure, we also indicated the data points at the characteristic frequency of ${\rm yr^{-1}}$ with the presupposed $-\frac{2}{3}$ power-law spectrum from the latest NANOGrav(Agazie et al., 2023a), EPTA(EPTA Collaboration et al., 2023) and CPTA(Xu et al., 2023) results. From the figure we can see that, our results generally exhibit the characteristic $-\frac{2}{3}$ power-law dependence. In the high-frequency region, the curve drops slightly, which can be easily understood as the frequency upper limit $f_{\rm ISCO}$ is different from event to event, according to Eq. (5), only events with small total mass can reach a relatively high signal frequency. This will lead to a decrease in the number of events in the high-frequency bins and a deviation from the $-2/3$ power-law behavior in the characteristic strain spectrum. On the other hand, it is easily noted that our curve lies below the observational results which have $A_{\rm yr^{-1}}\approx 4.85\times 10^{-16}$ at $f=1{\rm yr}^{-1}$. We will discuss this point in the discussiong section. Nevertheless, the projected value resides well within the theoretical range [$1\times 10^{-16}$, $5\times 10^{-15}$], as supported by existing literature.
Figure 6: The characteristic strain spectrum of GWB signal from our light cone mock data (green points). The latest results from NANOGrav (Agazie et al., 2023a), EPTA (EPTA Collaboration et al., 2023) and CPTA (Xu et al., 2023) at the characteristic frequency of $\rm yr^{-1}$ are also shown.
@@ 3.3 Anisotropy
Now we study the anisotropy properties of the GWB signal we get from our mock data, which is a result of GW sources and the GW strains come from them not being distributed isotropically in the virtual observable universe. The method we use allow us to estimate the anisotropy of GWB in a simulation environment more reflective to the real universe, as the cosmic large scale structure is incoporated. In this subsection, we will present how the sources in our simulation data are distributed across the observable universe and demonstrate the resulting anisotropic characteristics of the GWB signal.
3.3.1 Spatial Distribution of GW Sources
As mentioned previously, we obtain $N_{\rm total}\approx 3.027\times 10^{7}$ SMBBH GW sources in total. Due to the exceedingly large number, it is impractical to represent all of them in the figure. Thus we show the $\frac{1}{8}$-cube three dimensional spatial distribution of GW sources with side length $500$ Mpc in Figure 7, and its two dimensional projection in Figure 8. It can be seen from the figures that the supermassive black holes we constructed extended from the earth position to comoving distance beyond 8000 Mpc, corresponding to redshift from 0 to over 6. Similar patterns to the “cosmic web” can be recognized from the figures, demonstrating that supermassive black holes may also carry information about the large-scale structure of the universeShao et al. (2022); Zheng et al. (2023); Li et al. (2024), we will discuss this point in subsection 3.3.3.
Figure 7: The three dimensional spatial distribution of GW sources whose locations are within the box of side length 500Mpc ($0<x_{s},y_{s},z_{s}<500\rm Mpc$). Here we only show 1/8 part of the observable universe we constructed. We use colors to represent the GW amplitude of each source, we also mark the Earth’s position in the figure with a large blue sphere. From the figure we can distinguish the distribution patterns of cosmic large-scale structures such as voids and clusterings to a certain degree.
Figure 8: Left: The two dimensional projection of the three dimensional spatial distribution shown in Figure 7. Here we use colors to represent the coordinate $z_{s}$ of the GW sources. We can also distinguish void regions and fibrous structures from the figure.
@@ 3.3.2 Angular Power Spectrum
Figure 9: The sky map of GWB strain $h(\hat{\Omega})$ across the celestial sphere. Since energy density is proportional to $h^{2}(\hat{\Omega})$, it will follow similar sky map with each value on the sphere squared.
We transform the three dimensional Cartesian coordinates $(x_{s},y_{s},z_{s})$ of the mock SMBBH locations to spherical polar coordinates $(r,\theta,\phi)$, where $r=\sqrt{x_{s}^{2}+y_{s}^{2}+z_{s}^{2}}$, $\phi$ represents right ascension (the azimuthal angle respective to $x_{s}$-axis positive direction), and $\theta$ represents declination (elevating angle respective to $x_{s}Oy_{s}$ plane). By superimposing the gravitational wave signals from all possible sources along a particular direction, we obtain a background gravitational wave signal sky map distributed across the celestial sphere. In Figure 9, we show the resulted sky map for GW strain $h(\hat{\Omega})$. Since energy density $\rho(\hat{\Omega})$ is proportional to the squared strain $h^{2}(\hat{\Omega})$, the sky map of energy density will be similar to the strain map, with the value of each pixel squared.
Similar to Yang et al. (2019); Taylor & Gair (2013), we use spherical harmonic analysis to quantitatively analyze the anisotropy properties of the GWB signal from mock SMBBHs. The GW energy density $\rho(\hat{\Omega})$ can be expanded under the spherical harmonic basis:
@@
In practical, we divided the celestial sphere into $N$ small regions to perform the superposition of GW signals in a specific direction, so the integral above should be replaced with a summation:
@@
and the estimated error of each coefficient $C_{l}$ from cosmic variance is (Bartolo et al., 2022):
$\Delta C_{l}=\sqrt{\frac{2}{2l+1}}C_{l}.$ (9)
To calculate the coefficients $C_{l}$ for GW strain $h$, we just substitute energy density $\rho(\hat{\Omega})$ with strain $h(\hat{\Omega})$ in Equation (6), (7) and (8).
Based on these formulas, we calculated the angular power spectrum for the GWB energy density $\rho(\hat{\Omega})$. In Figure 10, we show the resulted $C_{l}$’s as a function of $l$. We can see that, $C_{1}/C_{0}\approx 2.50\times 10^{-3}\pm 2.04\times 10^{-3}$ and $C_{2}/C_{0}\approx 2.48\times 10^{-3}\pm 1.57\times 10^{-3}$, and higher-order $C_{l}/C_{0}$ values generally fluctuate within the range of 0.001 to 0.003. The behavior of the $C_{l}$ coefficients from energy density resides well within the observational constraints, i.e., $C_{l>0}/C_{l=0}<20\%$ (Agazie et al., 2023c). Fluctuations in the ratio $C_{l}/C_{0}$ across different values of $l$ may suggest the presence of clustering patterns in the spatial scale correspond to each specific $l$, and may indicate the existence of underlying structures in the angular distribution of the background GW strains. In Figure 11, we also plotted the $C_{l}$ coefficients in different frequency bins as shown in its legend. It can be seen from the figure that the ratio $C_{l}/C_{0}$ differ significantly across different frequency bins, and general increase with increasing frequency, this can be understood because as the frequency increases, the number of SMBBHs decreases, leading to an increase in anisotropy. $C_{1}/C_{0}$ at higher frequency end can even reach $\sim 0.8$. When $l\gg 1$, $C_{l}/C_{0}\simeq 0.6$ at high frequency end around $\sim 10^{-6}$ Hz. While in the lower frequency end, $C_{l}/C_{0}$ ranges from $0.001$ to $0.01$ for $l>0$. The total anisotropy mainly depends on lower frequency end around $\sim 10^{-9}$ Hz.
Figure 10: The angular power spectrum coefficients $C_{l}$ as a function of $l$ for GWB energy density $\rho$, where error bar indicate the cosmic variance.
Figure 11: The angular power spectrum coefficients $C_{l}$ as a function of $l$ for GWB energy density $\rho$ in different frequency bins.
3.3.3 GW Sources in a Shell

### v3

@@ 3 Results
We superimposed the gravitational wave signals from all the PTA band light-cone SMBBH events to obtain the GWB signal and studied its properties. In this section, we will present the basic characteristics of the SMBBH mock data we have obtained, as well as our main results for the GWB signal.
3.1 Statistical Properties of Light-Cone SMBBH Mock Data
We present the basic statistical properties of our light-cone SMBBH mock data in this subsection. The redshift $z$, observed frequency $f_{obs}$, total mass $M$ and mass ratio $q$ distributions of the SMBBHs of all the three hardening timescales are demonstrated in Figure 3. The redshift distribution of the SMBBHs peaks around redshift $z\sim 0.7$ and $z\sim 0.3$, and the maximum redshift can reach $z=6$ and $z=1$ for hardening timescale of $0.1\,\mathrm{Gyr}$ and $5\,\mathrm{Gyr}$ respectively. While for the $\tau_{h}=10\,\mathrm{Gyr}$ case, the majority of SMBBH events are distributed in the lowest redshift bin ($z<0.21$) [^5] In this extreme scenario, our data indicate that approximately 0.1% of galaxies with stellar mass $M_{\star}>10^{10}\,M_{\odot}$ below redshift z=0.21 host PTA-band SMBBHs. If it is possible for future experiments to conclusively identify nearby sub-pc SMBBHs, this occupation fraction may provide an additional observational test of our model., with the highest redshift not exceeding 0.5. The frequency distribution basically follow the power law distribution while subjecting to larger fluactuations in the higher frequency bins. The total masses of the binary systems peak around $3\times 10^{7}M_{\odot}$ for all the three hardening timescales. On the other hand, the mass ratio distribution of the SMBBHs generally exhibits a monotonically decreasing trend with increasing mass ratio $q$, which indicates that in our data, the majority of the binary black hole systems exhibit a significant difference in mass between the two black holes. Note that our assumption of a uniform hardening timescale may bias the mass ratio distribution. Crucially, binaries with comparable masses (high mass ratios) may experience shorter dynamical friction timescales and transition more rapidly into the GW-driven regime. This will result in a higher population of such systems evolving within the PTA frequency band, thereby increasing the number of high mass ratio sources captured in our observable light cone. The modeling of more refined hardening process is beyond the scope of this study, we’ll explore related issues in future work.
Figure 3: The redshift $z$ (upprt left), observed frequency $f_{\rm obs}$ (upper right), total mass $M$ (lower left) and mass ratio $q$ (lower right) distribution of the light-cone SMBBHs from one representative realization for all the three hardening timescales. The vertical black dashed lines in the frequency distribution plot indicate the lower frequency limits detectable with total observation durations of 10 years and 20 years respectively.
@@ 3.2 Isotropic Part of the Simulated GWB Signal
Our methodology incorporates a vast amount of individual sources scattered across the observable universe to construct GWB. For the sake of operational convenience, we considered the time interval for SMBBHs to evolve to coalescence when selecting light-cone events. Note that we do not account for binaries with frequencies above $f_{\rm max}=1/\Delta t$, as they lie beyond the observational window of PTA experiments under the parameter configuration we assumed. That is to say, we will impose a frequency window and only analyze the behavior of the resulting GWB signal between $f_{min}=1/T_{\rm obs}$ to $f_{max}=1/\Delta t$, where $T_{\rm obs}$ is the total observation duration and $\Delta t$ is the observation cadence.
@@
To calculate the resulting monopole part of the GWB signal and its characteristic strain spectrum, we divide the frequency range within the PTA frequency band spanning from $f_{\rm min}$ to $f_{\rm max}$ into different frequency bins. The bin width is determined by $\Delta f=1/T_{\rm obs}$, where $T_{\rm obs}$ is set to 10 years in this subsection. We select those SMBBHs whose observed frequencies are within the specified range in each bin. The resulting GW characteristic strain near a certain frequency can be calculated as a summation of strains from selected GW sources as:
$h_{\rm c}(f_{i})=\sqrt{\sum_{f_{i}<f_{j}\leq f_{i+1}}h_{j}^{2}f_{j}T_{\rm obs}},$
@@
where $h_{j}$ can be computed according to Equation (2), the summation over $j$ denote the superposition over all the GW events whose observed frequency $f_{j}$ falls within a specific frequency bin. We present our results regarding the characteristic strain spectrum of GWB for all the three hardening time scales in Figure 4. In this figure, we plot the results of all 50 realizations with light blue lines and show the average result of all realizations with a dark blue line. We also indicated the data points at the characteristic frequency of ${\rm yr^{-1}}$ with the presupposed $-\frac{2}{3}$ power-law spectrum from the latest NANOGrav (Agazie et al., 2023a), EPTA (EPTA Collaboration et al., 2023) and CPTA (Xu et al., 2023) results. From the figure, it is evident that as the hardening time increases, the averaged amplitude of GWB is suppressed, particularly when the hardening time reaches a scale of $10\,\mathrm{Gyr}$, where a significant reduction in the amplitude is observed. The characteristic amplitude $h_{c}$ at frequency $f=1\,\mathrm{yr}^{-1}$ is $4.08\times 10^{-16}$, $3.38\times 10^{-16}$, and $2.19\times 10^{-18}$ for $\tau_{H}=0.1\,\mathrm{Gyr}$, $5\,\mathrm{Gyr}$, and $10\,\mathrm{Gyr}$, respectively; while at $f=0.1\,\mathrm{yr}^{-1}$, $h_{c}=1.96\times 10^{-15}$, $1.42\times 10^{-15}$, and $3.13\times 10^{-16}$ for the same $\tau_{H}$ values in corresponding order. For the two cases with $\tau_{H}=$ 0.1 Gyr and 5 Gyr, where there are enough SMBBHs to demonstrate statistical trends, our results generally exhibit the characteristic $-\frac{2}{3}$ power-law dependence, although the higher probability of encountering bright sources in the low-frequency range has led to slight deviations in the averaged curves. On the other hand, in the high-frequency region the curve drops, which can be easily understood as the frequency upper limit $f_{\rm ISCO}$ is different from event to event, only events with small total mass can reach a relatively high signal frequency. This will lead to a decrease in the number of events in the high-frequency bins and a deviation from the $-2/3$ power-law behavior, the same reason has also led to the relatively large fluctuation observed in the higher frequency band of the characteristic strain spectrum. It can be noted that while the averaged curve lies below the observational results which have $A_{\rm yr^{-1}}\approx 2\times 10^{-15}$ at $f=1\,{\rm yr}^{-1}$, some random realizations can still reach observation levels at the characteristic frequency. Nevertheless, the projected averaged value resides well within the theoretical range [$1\times 10^{-16}$, $5\times 10^{-15}$], as supported by existing literature. We will discuss this point in the discussion section.
We also investigated whether resolvable individual sources exist in SMBBH samples under different hardening times for current PTA detectors. We assumed 68 pulsars, a 10-year observation time, and a timing residuals of 350ns, which are typical for current IPTA. We calculated the SNR for all SMBBH events in the sample and counted the number of events with SNR exceeding 8 or 3. The results are shown in Figure 5666The event with the largest SNR$>$200 in realization 21 has a distance of 0.67 Mpc, and the masses of the two SMBHs are $5.3\times 10^{9}\,M_{\odot}$ and $6.7\times 10^{7}\,M_{\odot}$ respectively. It coincidentally has both a relatively close distance and large black hole masses. We have checked that no SMBBH events come from the galaxy where Earth is placed.. It can be seen from the figure that generally the number of bright sources shows an inverse correlation with the hardening timescale. Considering that current PTA experiments haven’t detected any individual sources yet, we will subtract these bright individual sources when investigating the anisotropic properties of background signals and resovable individual sources under future PTAs to ensure that our results better matches the actual observations.
Figure 4: The characteristic strain spectrum of GWB signal from our light-cone SMBBH events. The three panels from left to right represent the results for hardening times of $0.1,5$ and $10\,\mathrm{Gyr}$ respectively. The light blue lines represent the results of different realizations, while the dark blue line indicates the average of all realizations. The latest results from NANOGrav (Agazie et al., 2023a), EPTA (EPTA Collaboration et al., 2023) and CPTA (Xu et al., 2023) at the characteristic frequency of $\rm yr^{-1}$ are presented in the figure as a comparison. The sensitivity curves of the three PTA configurations that we utilized for SNR calculations, including the current IPTA in this section and the future SKA-PTA and CPTA presented in Table 1 of Section 3.4, are also shown.
Figure 5: Left panel: The distribution of number of realizations versus different number counts of SMBBH events with SNR exceeding 8. Middle panel: The distribution of number of realizations versus different number counts of SMBBH events with SNR exceeding 3. Right panel: The maximum SNR in different realizations containing bright sources.
@@ 3.3 Anisotropy
Now we study the anisotropy properties of the GWB signal we get from our mock data, which is a result of GW sources and the GW strains come from them not being distributed isotropically in the virtual observable universe. The method we use allow us to estimate the anisotropy of GWB in a simulation environment more reflective to the real universe, as the cosmic large scale structure is incoporated. In this subsection, we will present the angular power spectrum of the GWB signal and the spatial distribution properties of the GW sources. Given the observational challenges associated with detecting anisotropy, a 20-year total observation duration was adopted here for characterizing anisotropy properties.
@@ 3.3.1 Angular Power Spectrum
Figure 6: The skymap of GWB relative energy density from one representative realization, where $\Omega_{GWB}$ is the ratio of GWB energy density to the critical density. The three panels from up to down represent the results for hardening times of $0.1,5$ and $10\,\mathrm{Gyr}$ respectively. The dark gray areas in the lowest panel denote the sky regions where no GW events are distributed.
We transform the three dimensional Cartesian coordinates $(x_{s},y_{s},z_{s})$ of the mock SMBBH locations to spherical polar coordinates $(r,\theta,\phi)$, where $r=\sqrt{x_{s}^{2}+y_{s}^{2}+z_{s}^{2}}$, $\phi$ represents right ascension (the azimuthal angle respective to $x_{s}$-axis positive direction), and $\theta$ represents declination (elevating angle respective to $x_{s}Oy_{s}$ plane). By superimposing the gravitational wave signals from all possible sources along a particular direction, sky maps of the background gravitational wave signal across the celestial sphere could be obtained. Similar to Yang et al. (2019); Taylor & Gair (2013), spherical harmonic analysis could be used to quantitatively analyze the anisotropy properties of the GWB signal from mock SMBBHs. The GW energy density $\rho(\hat{\Omega})$:
$\rho(\hat{\Omega})=\frac{\pi}{4G}\int f^{2}h^{2}_{c}(f,\Omega)d\ln f$ (3)
is the superposition of the energy densities across all frequency bins in a specific direction, and can be expanded under the spherical harmonic basis:
@@
In practice, the celestial sphere is divided into $N$ small regions to perform the superposition of GW signals in a specific direction, so the integral above should be replaced with a summation:
@@
We utilize Healpix to generate sky maps and calculate the corresponding angular power spectra of the GWB total energy density for all realizations. In Figure 6, we show the resulted sky maps for one representative realization for each of the hardening time scale. For the angular power spectrum, considering individual sources hasn’t been detected by current PTA experiments yet, we calculated the $C_{l}$ coefficients after bright sources (with SNR$>$8 or SNR$>$3) were subtracted and presented the SNR$<$8 case in Figure 7. We can observe from all the panels that $C_{l}/C_{0}$ exhibits significant fluctuations in all cases. Overall, for most of the realizations of $0.1$, $5$ and $10\,\mathrm{Gyr}$ hardening timescales, the distribution of $C_{1}/C_{0}$ falls within the range of $[3.0\times 10^{-2},4.3\times 10^{-1}]$, $[5.3\times 10^{-2},6.8\times 10^{-1}]$, $[7.7\times 10^{-2},6.2\times 10^{-1}]$ respectively (16%-84% distribution range). We also examined whether the level of anisotropy is sensitive to the substraction of additional bright sources by plotting in the left panel of Figure 8 the distribution of number of realizations versus the value of $C_{1}/C_{0}$ after bright sources with SNR$>$8 or SNR$>$3 were subtracted. From this figure we can see that, although cases with shorter hardening time scales tend to exhibit higher $C_{1}/C_{0}$ ratios, the distribution ranges of $C_{1}/C_{0}$ values for all three hardening timescales show significant overlap in general. While the removal of additional bright individual sources show negligible effect when $\tau_{H}=10\,\mathrm{Gyr}$ (This is expected, as only one realization contains a single source with SNR$>$3 in this case), it leads to a slight reduction in $C_{1}/C_{0}$ values for a small subset of realizations in the $\tau_{H}=0.1\,\mathrm{Gyr}$ and $\tau_{H}=5\,\mathrm{Gyr}$ cases. Collectively, the $C_{1}/C_{0}$ ratio of total energy density is insensitivity to the subtraction of extra bright sources. In Figure 9, we also plotted the $C_{1}/C_{0}$ value of energy density from all SNR$<$8 events across different frequency bins for all realizations. It can be seen from the figure that $C_{1}/C_{0}$ general increase with increasing frequency, this can be easily understood because as frequency increases, the number of SMBBHs decreases, leading to an increase in anisotropy. In the lowest frequency bin, $C_{1}/C_{0}$ falls mostly within the range of $[2.2\times 10^{-3},3.5\times 10^{-2}$], $[4.0\times 10^{-3},6.0\times 10^{-2}]$, and $[2.1\times 10^{-2},2.9\times 10^{-1}]$ for $\tau_{H}=0.1$, $5$, and $10\,\mathrm{Gyr}$ respectively (16%-84% distribution range). After substracting more bright sources with SNR$>$3, these distribution range further decreased to $[1.2\times 10^{-3},9.1\times 10^{-3}$], $[2.7\times 10^{-3},2.0\times 10^{-2}]$, and $[2.1\times 10^{-2},2.8\times 10^{-1}]$ for the same $\tau_{H}$ values in corresponding order. Compared to the total energy density distribution, all three hardening timescale cases exhibit lower anisotropy levels for energy density in the lowest frequency bin, and also demonstrating more divergent distribution ranges among themselves. We have checked that this arises because low-frequency events dominate the SMBBH population, while the $f^{2}$ factor in the energy density integral in Eq. (3) does not affect the $C_{1}/C_{0}$ ratio within individual frequency bins, it does modify this ratio for the total energy density by preferentially weighting higher-frequency contributions. Thus the directional dependence of GW frequency distributions become an important source of anisotropy in the total energy case, which introduces extra anisotropy in the distribution, and also mitigates the differences between results of different hardening timescales. As a comparison, we also plotted the distribution of number of realizations versus the value of $C_{1}/C_{0}$ for energy density in the lowest frequency bin after bright sources with SNR$>$8 or SNR$>$3 were subtracted in the right panel of Figure 8. As clearly demonstrated in the figure, the subtraction of additional bright sources yields greater reduction in anisotropy for energy density in the lowest frequency bin compared to the total energy density case, indicating that bright individual sources has more important contribution to anisotropy in this case.
The distribution range of the power spectrum coefficients and their frequency-dependence are generally consistent with the results of Gardiner et al. (2024), while our results demonstrate relatively lower anisotropy in the lowest frequency bin compared to this literature (which may be attributed to the fact that we have subtracted unrealistic bright sources that haven’t been detected by current PTAs yet), and the fluctuations appear more pronounced in higher frequency bins (which may be attributed to the incorporation of large-scale structure in our sample distribution). It also worth mention that in the lowest frequency bin, the behavior of the $C_{l}$ coefficients from most realizations resides well within the observational constraints, i.e., $C_{l=1}/C_{l=0}<20\%$ (Agazie et al., 2023c), however, in the high-frequency range, this limit is more likely to be exceeded, a manifestation that aligns with the results reported in Gardiner et al. (2024).
Figure 7: The angular power spectrum coefficients $C_{l}$ as a function of $l$ for GWB energy density $\rho$ after subtracting the events with SNR greater than 8 for all the 50 realizations of each hardening timescale. Panels from left to right represent the results for hardening times of $0.1,5,10\,\mathrm{Gyr}$ respectively.
Figure 8: Left panel: the distribution of number of realizations versus the value of $C_{1}/C_{0}$ of total energy density for all the three hardening timescales. Right panel: the distribution of number of realizations versus the value of $C_{1}/C_{0}$ for energy density in the lowest frequency bin for all the three hardening timescales. In both panels solid lines represent results after subtracting the events with SNR greater than 8, and dashed lines represent results after subtracting the events with SNR greater than 3. In the right panel, for $\tau_{H}=10$ Gyr case, since bright sources are absent in the lowest-frequency bin, the green line is completely unaffected by subtraction of additional bright sources.
Figure 9: The angular power spectrum coefficients $C_{1}/C_{0}$ of GWB energy density $\rho$ in different frequency bins for all the three hardening time scales. The three panels from left to right represent the results for hardening times of $0.1,5,10\,\mathrm{Gyr}$ respectively. The light blue lines represent the results of different realizations, while the red lines indicate the average of all realizations.
3.3.2 GW Sources in Shells